OpenStudy (anonymous):

Whats the y intercept of the line that passes through the point (3,-7) and (-1,9) woos CooperP -.-

OpenStudy (anonymous):

<__>

OpenStudy (anonymous):

you'll need an equation to solve y y=mx+c to find the equation, you'll need the gradient/slope, which is 'm' on the equation above so  your gradient is -4 and so y=-4x+c to find the c, which is your y-intercept, simply substitute in the x and y from either points i'll go with the (3,-7) one y=mx+c y=-4x+c -7=-4(3)+c -7+4(3)=c can you solve for c (y intercept)?